Which of the following real valued functions is/are not even functions?

(A)

f(x) = x3 sin x
(B) f(x) = x2cos x
(C) f(x) = exx3 sin x
(D) f(x) = x–[x], where [x] denotes the greatest integer less than or equal to x

Grade:

1 Answers

Vikas TU
3 years ago
Test to determine if a function y=f(x) is even, odd or neither: Replace x with -x and compare the result to f(x). If f(-x) = f(x), the function is even. If f(-x) = - f(x), the function is odd.